There are three naturally occurring isotopes
of the hypothetical element hilarium 40Hi,
41Hi, and 43Hi. The percentages of these
isotopes are the following:
Hilarium Natural
Isotopes Abundance
40Hi 29.4%
41Hi 37.3%
43Hi 33.3%
You can assume that the atomic masses are
exactly equal to the mass number. Calculate
the weight of “naturally” occurring hilarium
and report it as you would for any other naturally occurring element.
1
Expert's answer
2020-10-30T06:45:33-0400

Q141268

There are three naturally occurring isotopes of the hypothetical element hilarium 40Hi,

41Hi, and 43Hi. The percentages of these isotopes are the following:

Hilarium Isotopes Natural Abundance in %

40 Hi 29.4%

41 Hi 37.3%

43Hi 33.3%

You can assume that the atomic masses are exactly equal to the mass number. Calculate

the weight of “naturally” occurring hilarium and report it as you would for any other naturally occurring element.


Solution :  


We can use the formula of Average atomic mass for finding atomic mass of Hilarium.



Average atomic mass = Σ (mass # of Isotope * natural abundance ) / 100



Average atomic mass = ( 40 * 29.4 + 41 * 37.3+ 43 * 0.333) / 100


= (1176 + 1529.3 + 1431.9)/100

= 41372/100

= 41.372 amu



in 3 significant figure the answer is 41.4 amu or 41.4 g/mol .


Please also refer to the images I have attached with this for more explanation.